OpenStudy (anonymous):

evaluate the expression when a=3 4a^4 A. 324 B. 48 C. 28 D. 11

geerky42 (geerky42):

plug in 3. 4(3)^4

geerky42 (geerky42):

Can you calculate 4(3)^4?

OpenStudy (anonymous):

so that would be... 4*3=12.. 12^4 = 20, 736 im confused..

geerky42 (geerky42):

Remember PEDMAS? Parentheses first, but it is already done, so now exponents. Do exponents first. After this, next is multiplication and division. Multiply.

geerky42 (geerky42):

So you do 3^4 first then multiple it by 4.

OpenStudy (anonymous):

OHH YEA I DIDNT THINK ABOUT THAT. so its 3^4=81, 81*4=324 which is A. thank you.

geerky42 (geerky42):

Welcome!